# Microcanonical simulated annealing: Massively parallel Monte Carlo simulations with sporadic random-number generation

### v1

@@ Appendix A Computation of the best time shift
As we have explained in Sec. III, to estimate the best value of $k^{*}$, we use the coherence length $\xi_{12}$. A similar analysis can be performed with the energy, but, as it is clear from panel d in Fig 2, the energy in the microcanonical algorithm does not decrease to its equilibrium value monotonically. Instead, for short times, it is sensitive to the daemons’ refreshing process. This behavior makes it more challenging to estimate $k^{*}$. In addition, the relation between the coherence length and time in the spin-glass phase, which can be extended to the critical point, offers a direct way to estimate $k^{*}$.
A possible interpolating function is [42]
@@
at $T_{\mathrm{c}}$, where $\omega$ is the exponent characterizing the first corrections to leading scaling, $\omega=1.12(10)$ [39], while for $T<T_{\mathrm{c}}$, $\omega$ is the replicon exponent [42] (we used $\omega=0.35$ in the fits).
Then, by a simple time shift $k$, we obtain
$\log(t)=a_{0}-\log(k)+z\log(\xi_{12})+\frac{a_{1}}{\xi_{12}^{\omega}}\,.$ (25)
Our procedure starts by fitting the canonical data to Eq. (24) to obtain the parameters $a_{0}$, $a_{1}$, and $z$. Finally, we fit the microcanonical data, with $k$ as the only fitting parameter, which results in $k^{*}$. Following this process, we obtain the results summarized in Table 1.
The only exception is the value of $k^{*}$ at $T=1.4$. In the case of data in the paramagnetic phase, this approach is impractical because Eqs. (24) and (25) do not hold. In this case, we interpolate $\xi_{12}(t)$ for the canonical data to the following functional expression
$\xi_{12}(t)=C\left(1-\sum_{i=0}^{7}a_{i}\mathrm{e}^{-tb^{i}/\tau}\right)\,,$ (26)
(see the fit result in Table 2). Then, by taking $g(t)=\xi_{12}(kt)$ we can obtain the value of $k^{*}$ by fitting our microcanonical data with $k$ as the only fitting parameter.
@@ Appendix B The CUDA implementation
In the following, we provide basic technical information about the CUDA implementation of the mic.SA algorithm described in Sec. II applied to the three-dimensional Ising spin glass. The code implements a two-level parallelism: we use multispin coding to pack 64 spins in 64-bit words (long long int) and then CUDA threads for the concurrent update of even and odd spins. Multiple replicas and samples can be simulated in a single execution; the actual number of both depends on the resources (mainly memory) available on the GPU (we successfully tested up to 512 replicas). The code also implements the standard Metropolis algorithm using random numbers generated on the GPU with the procedure described in Ref. [25]. The refresh of the daemons/walkers is carried out on the CPU asynchronously to the microcanonical update (we recall that the refresh takes place according to an exponential schedule) using CPU threads. The walking procedure and the microcanonical update are implemented in two separate CUDA kernels to increase flexibility. The main reason is that the microcanonical update is carried out separately for each daemon/walker, whereas the walking procedure is completed for all the daemons in a single kernel execution.
Table 3: Metropolis and mic.SA timings for a full sweep of 128 replicas of a lattice with $160^{3}$ spins. The total time of the Metropolis sweep is 1.9 times longer than that of mic.SA.
@@ Acknowledgements.
This work was partially supported by MICIU/AEI/10.13039/501100011033 (Spain) and by “ERDF: A way of making Europe” (European Union) through grants no. PID2022-136374NB-C21, PID2020-112936GB-I00 and RED2022-134244-T; by Junta de Extremadura (Spain) through grant no. GR24022; and by funding from the 2021 first FIS (Fondo Italiano per la Scienza) funding scheme (FIS783 - SMaC - Statistical Mechanics and Complexity) from Italian MUR (Ministry of University and Research). We acknowledge the use of the CESAR computational resources at the BIFI Institute (University of Zaragoza). IG, MB and FRT acknowledge the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 – CUP E83C22003230001, CUP B93C22000620006 and CUP B83C22002940006, Mission 4 Component 2 Investment 1.4, funded by the European Union – NextGenerationEU.

### v2

@@ Appendix A Computation of the best time shift
As we have explained in Sec. III, to estimate the best value of $k^{*}$, we use the coherence length $\xi_{12}$. A similar analysis can be performed with the energy, but, as it is clear from panel d) in Fig 2, the energy in the MicSA algorithm does not decrease to its equilibrium value monotonically. Instead, for short times, it is sensitive to the daemons’ refreshing process. This behavior makes it more challenging to estimate $k^{*}$ from these data. Nonetheless, the relation between the coherence length and time in the spin-glass phase, which can be extended to the critical point, offers a straightforward way to estimate $k^{*}$.
In particular, at the critical point, we can use the following relation [46]
@@
where $\omega$ is the exponent characterizing the first corrections to leading scaling and it takes the value $\omega=1.12(10)$ [43].
For $T<T_{\mathrm{c}}$, the leading behavior is given by $\omega=\theta$, where $\theta$ is the replicon exponent (we used $\omega=0.35$ in the fits [46]).
As the reader will notice, it is easy to include in Eq. (24) a time shift $k$ that connects our MicSA data with the canonical curve:
$\log(t)=a^{\prime}_{0}-\log(k)+z^{\prime}\log(\xi_{12})+\frac{a^{\prime}_{1}}{\xi_{12}^{\omega}}\,.$ (25)
In this situation, we propose the following method to extract the best time shift, $k^{*}$:
  1. 1.
    Fit the canonical data to Eq. (24) in order to obtain parameters $a_{0}$, $a_{1}$, and $z$.
  2. 2.
    With parameters fixed to the values obtained in the previous step ($a^{\prime}_{0}=a_{0}$, $a_{1}^{\prime}=a_{1}$, and $z^{\prime}=z$), we perform a new fit to the MicSA data to Eq. (25).
This process guarantees that the only fitting parameter in step 2 is $k$, obtaining $k^{*}$. Table 1 summarizes the results of $T\leq T_{\mathrm{c}}$. Fig. 5 reproduces this procedure. In the main panel, we report the difference between our MicSA data $\log(t_{\mathrm{MicSA}})$ and the interpolation $f(t;k^{*}|\Theta)$, see Eq. (25), with the best time shift $k^{*}=5.875$ and the parameters $\Theta=(a_{0},a_{1},z)$ obtained following this process for $\gamma=6$ and $T=T_{\mathrm{c}}$. The inset shows the non-subtracted data.
This approach is, however, not useful at $T>T_{\mathrm{c}}$, because Eq. (24) does not hold. In this regime, we interpolate $\xi_{12}(t)$ as a function of time using the following expression
$f(t)=C\left(1-\sum_{i=0}^{7}a_{i}\mathrm{e}^{-tb^{i}/\tau}\right)\,.$ (26)
In this situation, we can use a similar protocol to extract the value of $k^{*}$:
  1. 1.
    We fit our canonical data to Eq. (26) to obtain the values of the parameters (see the fit result in Table 2).
  2. 2.
    Using the parameters obtained in the previous step, we fit our MicSA data to $f(kt)$, where $k$ is the only fitting parameter.
As in the case of $T\leq T_{\mathrm{c}}$, this process allows us to obtain directly the best time shift $k^{*}$, even in the case of $T>T_{\mathrm{c}}$. Notice that, in this situation, $t$ is the independent variable. Then, in Table 1 we reported the fitting range in $t$, instead of $\xi_{12}$. Notice that, although we use a slightly different process to obtain the value of $k^{*}$ for $T>T_{\mathrm{c}}$, panels d and e in Fig. 4 reported a value of $\delta_{e}$ and $\delta_{\xi_{12}}$ compatible with zero at long times.
@@ Appendix B The CUDA implementation
In the following, we provide basic technical information about the CUDA implementation of the MicSA algorithm described in Sec. II applied to the three-dimensional Ising spin glass. The code implements a two-level parallelism: we use multispin coding to pack 64 spins in 64-bit words (long long int) and then CUDA threads for the concurrent update of even and odd spins. Multiple replicas and samples can be simulated in a single execution; the actual number of both depends on the resources (mainly memory) available on the GPU (we successfully tested up to 512 replicas). The code also implements the standard Metropolis algorithm using random numbers generated on the GPU with the procedure described in Ref. [25]. The refresh of the daemons/walkers is carried out on the CPU asynchronously to the microcanonical update (we recall that the refresh takes place according to an exponential schedule) using CPU threads. The walking procedure and the microcanonical update are implemented in two separate CUDA kernels to increase flexibility. The main reason is that the microcanonical update is carried out separately for each daemon/walker, whereas the walking procedure is completed for all the daemons in a single kernel execution.
Table 3: Metropolis and MicSA timings for a full sweep of 128 replicas of a lattice with $160^{3}$ spins. The total time of the Metropolis sweep is 1.9 times longer than that of MicSA.
@@ Acknowledgements.
We thank Martin Hasenbusch for calling our attention to Refs. [33, 34, 35, 36] and for kindly explaining to us some of the details in these works. This work was partially supported by MICIU/AEI/10.13039/501100011033 (Spain) and by “ERDF: A way of making Europe” (European Union) through grants no. PID2022-136374NB-C21, PID2020-112936GB-I00 and RED2022-134244-T; by Junta de Extremadura (Spain) through grant no. GR24022; and by funding from the 2021 first FIS (Fondo Italiano per la Scienza) funding scheme (FIS783 - SMaC - Statistical Mechanics and Complexity) from Italian MUR (Ministry of University and Research). We acknowledge the use of the CESAR computational resources at the BIFI Institute (University of Zaragoza). IG, MB and FRT acknowledge the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 – CUP E83C22003230001, CUP B93C22000620006 and CUP B83C22002940006, Mission 4 Component 2 Investment 1.4, funded by the European Union – NextGenerationEU.